### v1

# Spatial correlations in SIS processes on regular random graphs

Abstract
In network-based SIS models of infectious disease transmission, infection can only occur between directly connected individuals. This constraint naturally gives rise to spatial correlations between the states of neighboring nodes, as the infection status of connected individuals becomes interdependent. Although mean-field approximations are commonly invoked to simplify disease forecasting on networks, they fail to account for these correlations by assuming that infectious individuals are well-mixed within a population, leading to inaccurate predictions of infection numbers over time. As such, the development of mathematical frameworks that account for spatially correlated infections is of great interest, as they offer a compromise between accurate disease forecasting and analytic tractability. Here, we use existing corrections to mean-field theory on the regular lattice to construct a more general framework for equivalent corrections on regular random graph topologies. We derive and simulate a system of ordinary differential equations for the time evolution of the spatial correlation function at various geodesic distances on random networks, and use solutions to this hierarchy of ordinary differential equations to predict the global infection density as a function of time, finding good agreement with corresponding numerical simulations. Our results constitute a substantial development on existing corrections to mean-field theory for infectious individuals in SIS processes and provide an in-depth characterization of how structural randomness in networks affects the dynamical trajectories of infectious diseases on networks.
I Introduction
The Susceptible-Infected-Susceptible (SIS) model, first introduced by Kermack and McKendrick [1], is a foundational framework for describing the spread of infectious diseases in which individuals do not acquire lasting immunity following infection (e.g. the common cold or many sexually transmitted diseases [2, 3, 4, 5]). In the model, members of a population transition repeatedly between two states: susceptible (S) and infected (I). Many generalized SIS models assume that susceptible and infected populations are well-mixed spatially, such that the rates of contact between infected and susceptible individuals are assumed to be uniform across a population. Such models are often referred to as mean-field approximations, and, while they are highly amenable to mathematical analysis [6, 7, 4], they are generally inappropriate for real-world scenarios, in which infected individuals are spatially ordered or clustered. In such contexts, the dynamical trajectory of a disease is highly sensitive to the spatial distribution of infected individuals, where infection is spread more rapidly when infected individuals are distributed more evenly in space.
Figure 1: (a) Schematic of the SIS model on networks. Infected nodes recover with rate $\gamma$ (conventionally, time is rescaled as $t\to\gamma t$, implying $\gamma=1$) and infect their nearest neighbors, each at rate $\beta$. (b)-(e) Snapshots of SIS model simulations on RRGs in the quasi-steady state. For lower values of average node degree $k_{0}$, the clustering of infectious nodes (orange) is more pronounced than on networks with a higher node degree. Moreover, increasing the infection rate, $\beta$, also decreases the clustering of infectious nodes but increases the overall infection density in the steady state. (f) Steady state infection density on the RRG as a function of average node degree, $k_{0}$, for simulations on RRGs with $N=10^{4}$ nodes, $\beta k_{0}=1.7$, and $\gamma=1$. The dashed line is the expected mean-field value of $1-1/(\beta k_{0})\simeq 0.41$. The initial conditions of all simulations are an initial density of $x_{I}=10^{-2}$ with $100$ infective individuals placed randomly on the network, and all results are averaged over 100 simulations.
One solution to mitigate the inaccuracies of mean-field models is to represent the SIS process on a network, in which nodes represent individuals in a population and edges represent connections between individuals (that is, potential routes for the flow of infection). The dynamics of the SIS model on networks are governed by two processes – the infection per connection between infected and susceptible individuals, occurring at a rate $\beta$, and the recovery of infectious individuals at a rate $\gamma$ (see Fig. 1a).
A significant advance in the mathematical modeling of epidemic dynamics on networks was the development of pairwise approximation methods. In contrast to traditional mean-field models, pairwise models explicitly track the densities of connected node pairs of various types, capturing the essential local correlations that arise from network structure and disease transmission mechanisms. Such models established the formulation and analytic power of pairwise ODE frameworks, demonstrating their improved accuracy in regimes where local correlations are significant, such as networks with low degree or heterogeneous topologies [8, 9, 10].
Building on this foundation, modeling SIS processes on a network leads to the emergence of spatial correlations between infectious individuals, since transmission can only occur via direct connections to currently infected neighbors (see Figs. 1b-e). This local transmission mechanism leads to epidemic trajectories that differ markedly from those predicted by mean-field models, especially on networks with lower average node degree, where these deviations are more pronounced (Fig. 1f), or regular lattices. Existing corrections to mean-field theory offer an ordinary differential equation (ODE)-based approach for forecasting disease dynamics, relying on closure approximations that assume spatial correlations between infectious individuals decrease rapidly with shortest path distance [8, 11, 12]. Similarly, for networks with higher degrees of topological randomness—such as random perturbations on the regular lattice achieved through stochastic edge swapping algorithms—mean-field theory is also inaccurate; yet, no analytical framework exists to predict the spread of infection. Consequently, analyzing SIS processes on such networks remains challenging, and simulation-based approaches are typically preferred over analytical methods for predicting disease spread.
To address this problem, this work extends previous corrections to mean-field theory to derive a general framework for the accurate forecasting of disease spread in SIS processes on random network topologies. Motivated by numerical simulations that highlight the strong dependence of spatial correlations on underlying network structure, we first analyze spatial correlations in infection dynamics on a series of regular networks with increasingly randomized connectivity. Focusing specifically on the limiting case of the regular random graph (RRG), we then generalize existing lattice-based theory, deriving a system of ODEs to model the temporal evolution of spatial correlations at varying shortest path lengths. Unlike traditional pairwise models—which only account for correlations between directly connected (nearest-neighbor) nodes—our approach resolves the full hierarchy of spatial correlations at every network distance by solving coupled ODEs for each shell around an infected node. This enables us to capture how infection clustering decays with distance and how finite network connectivity and the network’s structure shape epidemic trajectories, which are missed by standard closures. Comparison with numerical simulations demonstrates that our derived methodology yields substantially improved accuracy over mean-field predictions and lattice-based corrections. Furthermore, by systematically refining mean-field theory with spatial correction terms that account for local clustering, we derive closed-form asymptotic expansions for both the endemic infection density and the two-point correlation function in the long-time limit. These steady-state approximations not only quantify the impact of network structure on endemic persistence, but also provide expressions that show how finite connectivity suppresses infection prevalence relative to classical mean-field predictions. Thus, our results offer a robust analytical tool for understanding and forecasting the dynamics of infectious diseases in SIS processes on RRGs, revealing how structural randomness affects the spatial organization of infections and thereby shapes epidemic trajectories.
II Corrections to SIS mean-field theory on regular lattices
The mean-field mathematical representation of SIS dynamics executed on the RRG makes the simplifying assumption that susceptible and infected populations are well-mixed spatially, permitting the construction of two simple ODEs governing the dynamics of the system
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\gamma x_{I}-\beta k_{0}x_{I}x_{S},\;\;\;\frac{\mathrm{d}x_{S}}{\mathrm{d}t}=\beta k_{0}x_{I}x_{S}-\gamma x_{I},$ (1)
where $\beta>0$ is the rate of infection of one susceptible individual from one infectious individual, $\gamma>0$ is the recovery rate of each infectious individual, $k_{0}$ is the fixed degree of each node, and $0\leq x_{I},x_{S}\leq 1$ are the population fractions in the $I$ and $S$ states, respectively. Here, the ratio $R_{0}=\beta k_{0}/\gamma$ is known as the basic reproduction number, a summary statistic that, given a mean-field assumption, determines whether an infection will eventually vanish from a population or become endemic [13].
The conservation law $x_{S}+x_{I}=1$ means that every individual is either susceptible or infected at any point in time, and time is often rescaled as $t\to\gamma t$ by setting $\gamma=1$. Hence, Eq. (1) can be reduced to a single equation
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{I}(1-x_{I})-x_{I}.$ (2)
In Eq. (2) the first term represents the infection of $k_{0}$ susceptible neighbors by an infectious node, each at a rate $\beta$, while the second term represents the recovery of infectious individuals at a rate of $1$. Here, $x_{I}(1-x_{I})$ gives the probability that a randomly selected node is susceptible and its neighbor is infected, assuming, under the mean-field approximation, that all nodes share the same average infection probability, independent of their neighborhood. Yet, due to the localized nature of the infection, it is more likely that a node will be infected if its neighbor is infected as well, which significantly affects the predicted trajectory of a disease for low $k_{0}$, leading to inaccurate forecasting (red and black traces in Fig. 2).
The same limitations of mean-field theory appear in network structures with non-random connectivity, such as lattice networks. For this reason, many applications require refining mean-field approaches to achieve more reliable forecasting. Previous extensions to mean-field theory, especially in the context of birth-death-movement models, have established a strong foundation for accurately predicting population dynamics on lattices by systematically including spatial correlations into the analysis [11, 12]. Here, each lattice site can be either occupied $A$ or empty $\emptyset$. The allowed processes include: birth, where a cell can reproduce into an adjacent empty site $A+\emptyset\to A+A$; death, where a cell is removed, leaving an empty site $A\to\emptyset$; and movement, where a cell moves to a neighboring empty site $A+\emptyset\to\emptyset+A$. Of these, the birth and death processes can be directly mapped to infection $I+S\to I+I$ and recovery $I\to S$ events in SIS dynamics, while cell movement is omitted. In this context, corrections to mean-field theory for SIS processes can be derived under the assumption that the positions and connectivity of individuals remain invariant over time.
The infection process in network-based SIS models relies on transmissions occurring exclusively across edges connecting susceptible and infected individuals. This local interaction structure makes it advantageous to formulate the dynamics directly in terms of the density of susceptible-infected pairs, $x_{SI}^{(1)}$, rather than only global densities [8]. Here and below, the superscript $1$ denotes distance-one adjacent lattice sites. Thus, the evolution equation for the infected fraction can be written as
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{SI}^{(1)}-x_{I}.$ (3)
To quantify deviations from mean-field predictions, we introduce pairwise correlation functions $F_{II}^{(1)}\!(t)$ and $F_{SI}^{(1)}\!(t)$, which measure the actual fractions of infected-infected and susceptible-infected pairs, compared to their mean-field values. More generally, for any network distance $r$, these are defined as
$F_{II}^{(r)}(t)=\frac{x_{II}^{(r)}(t)}{x_{I}(t)^{2}},\quad F_{SI}^{(r)}(t)=\frac{x_{SI}^{(r)}(t)}{x_{I}(t)(1-x_{I}(t))},$ (4)
where $x_{II}^{(r)}(t)$ and $x_{SI}^{(r)}(t)$ denote the fractions of infected-infected and susceptible-infected pairs at distance $r$, respectively. The parameter $r$ can refer to either Euclidean distance or shortest path length, but does not explicitly enter the system dynamics, as demonstrated in the following section. The normalization by mean-field edge densities, which assume statistical independence, ensures that $F_{II}^{(r)}$ and $F_{SI}^{(r)}$ quantify the degree of spatial correlation between node states at distance $r$ on the network.
Treating $x_{SI}$ and $x_{II}$ as edge probabilities, it follows that $x_{II}^{(r)}+x_{SI}^{(r)}=x_{I}^{(r)}=x_{I}$. This is since $x_{I}^{(r)}$ is the marginal probability of a node at distance $r$ from the reference node (regardless of its status) to be infected, and according to the mean-field assumption is equal to the fraction of infected individuals across the whole network. Therefore, the correlation functions are related as
$F_{SI}^{(r)}(t)=\frac{x_{I}(t)-x_{II}^{(r)}(t)}{x_{I}(t)(1-x_{I}(t))}=\frac{1-F_{II}^{(r)}(t)x_{I}(t)}{1-x_{I}(t)}.$ (5)
Using Eqs. (4) and (5) in Eq. (3), the pairwise model becomes an explicit correction to the mean-field equation,
$\frac{\mathrm{d}x_{I}}{\mathrm{d}t}=\beta k_{0}x_{I}\left[1-F_{II}^{(1)}\left(t\right)x_{I}\right]-x_{I}.$ (6)
This form illustrates how local clustering of infection, as captured by $F_{II}^{(1)}$, alters the epidemic trajectories predicted by mean-field theory.
The infected-infected pairwise correlation function $F_{II}^{(r)}\left(t\right)$ approaches unity when the infection states of nodes at distance $r$ are statistically independent, which occurs in well-mixed networks ($k_{0}\gg 1$), where infected nodes are expected to be randomly distributed without pronounced spatial clustering. Conversely, $F_{II}^{(r)}\left(t\right)$ is greater than unity when infected nodes exhibit clustering (positive spatial correlation), as commonly seen in network-based SIS processes due to local infection spread from neighbors, making infected nodes more likely to be found near each other than by chance. Under typical SIS dynamics on connected networks, values of $F_{II}^{(r)}\left(t\right)<1$ do not occur, since infection propagation inherently generates clustering rather than dispersion. Therefore, the term $1-F_{II}^{(1)}\!(t)x_{I}\;$ may be interpreted as an effective susceptible density, and is typically lower compared to the naive mean-field value. Similarly, the behavior of $F_{SI}^{(r)}(t)$ tends to lie below unity at intermediate distances due to local depletion of susceptible neighbors around infected sites. This attenuation reflects the saturation of infection fronts and limits the effective reach of contagion, beyond what the mean-field susceptible density would suggest.
Figure 2: Average infection density, $x_{I}$ (top row), and the distance-one pairwise correlation function, $F_{II}^{(1)}\left(t\right)$ (bottom row). Here, SIS model simulations on the RRG (red) are compared with numerical solutions of the corrections to mean-field theory derived in Section V (blue), and with mean-field theory (black). Simulations were executed on networks with $N=10^{4}$ nodes, with an initial infection density of $10^{-2}$ such that 100 infective individuals were placed randomly on the network, and with $\gamma=1$. The first two columns correspond to a node degree of $k_{0}=4$ with infection rates of $\beta k_{0}=1.7$ and $2.0$, while the last two columns correspond to $k_{0}=6$ with $\beta k_{0}=1.4$ and $1.7$, respectively. Results in each plot are averaged over $100$ simulations.
III Dependence of nearest-neighbor correlation on network topology
We now further motivate the development of bespoke corrections to mean-field theory for the RRG by using numerical simulations to highlight significant variations in $F_{II}^{(1)}$, and hence global infection dynamics, with the underlying topology of regular networks. In order to do this, we investigate the effects of rewiring the regular lattice in two and three dimensions by randomly exchanging edges on the network with probability $p$, inspired by the methodology of Watts and Strogatz [14]. Here, we construct networks that preserve the degree of nodes, but become increasingly random in structure, in the sense that the probability of two adjacent nodes being connected by an edge decreases as $p$ increases. The edge swapping algorithm we use is a modification of an algorithm from Wang et al. [15] and can be summarized as follows:
  • Consider a network $G(V,E)$ representing a regular lattice, where $E$ and $V$ are the edge and node sets.
  • Sample edges from $E$ randomly without replacement to create a list of all edges in $E$ in random order, say $L=(e_{1},e_{2},\dots,e_{|E|})$.
  • Select a variable $p\in[0,1]$, representing the probability of edge exchange on the lattice.
  • For each consecutive edge $(e_{2i-1},e_{2i})$ in $L$, sample a random number $p_{e}\sim U(0,1)$.
  • If $p_{e}\leq p$ swap the edges $e_{2i-1}=(u_{2i-1},v_{2i-1})$ and $e_{2i}=(u_{2i},v_{2i})$ to $e_{2i-1}^{\prime}=(u_{2i-1},v_{2i})$ and $e_{2i}^{\prime}=(u_{2i},v_{2i-1})$; Otherwise, keep the edges as they are, $e_{2i-1}^{\prime}=e_{2i-1}$ and $e_{2i}^{\prime}=e_{2i}$.
  • The revised network is $G=(V,E^{\prime})$ with $E^{\prime}=\{e_{1}^{\prime},e_{2}^{\prime},\dots,e_{|E|}^{\prime}\}$.
This algorithm yields randomized network structures that converge to the RRG as $p$ becomes ${\cal O}(1)$.
In studying how increasingly random connectivity on the network affects spatial correlations between adjacent nodes, we first conduct a parameter sweep over the total rate of infection for each node, $k_{0}\beta$, and the probability of edge swapping, $p$, on the two- and three-dimensional square lattices, for which the node degree is $k_{0}=4$ and $k_{0}=6$, respectively. Varying $k_{0}\beta$ and $p$, we find that both an increase in $p$ (and hence, a reduced mean shortest path length on the network) and a higher rate of infection per node, $k_{0}\beta$, suppress spatial correlations between infectious nodes. This is manifested in shorter relaxation times (approximated by the time between peak and half-maximum values of $F_{II}^{(1)}(t)$) for lower mean-shortest path lengths and higher rates of infection (Figs. 3(a) and 3(c)). Furthermore, the steady-state value of $F_{II}^{(1)}\left(t\right)$ (Figs. 3(b) and 3(d)), which is in turn related to the steady-state infection density on the network, is also highly sensitive to both $p$ and the rate of infection. Similar effects have previously been observed on the proportion of recovered individuals in the closely-related SIR model [16, 17].
Figure 3: Relaxation time and steady state values for the SIS process ($\gamma=1,N=10^{4},I_{0}=10^{-2}$) versus $k_{0}\beta$ and $p$. Figures (a) and (c) show log relaxation times, which we define as the time between the peak ($t_{\rm{peak}}$) and half-maximum ($t_{1/2}$) values of $F_{II}^{(1)}(t)$, whilst (b) and (d) show the value of $\ln(\bar{F}_{II}^{(1)}-1)$ for $k_{0}=4$ and $6$, respectively. The results of the parameter sweep show that increasing $p$ and increasing $k_{0}\beta$ both reduce the steady state value of $F_{II}^{(1)}\left(t\right)$ and the relaxation time for the process.
From the results of the parameter sweep, we see that changes in both $p$ and $k_{0}\beta$ strongly affect $F_{II}^{(1)}\left(t\right)$, and hence, the trajectory of infection density on the network as a function of time (via Eq. (6)). In general, adding topological randomness to the network reduces the relaxation time of the process and increases the steady-state infection value. As such, it is analogous to increasing $\beta$ in terms of its effect on spatial correlations between infectious nodes. The results of these simulations suggest that the corrections to mean-field theory proposed and analyzed in previous work on the regular lattice, where $p=0$ [11, 12], are inappropriate corrections for equivalent simulations on the RRG, where $p={\cal O}(1)$, due to a strong relationship between $p$ and $F_{II}^{(1)}\left(t\right)$. As such, in the proceeding section, we derive a novel correction to mean-field theory based on the theory of random networks, finding excellent agreement with corresponding numerical simulations that outperform both mean-field theory and previous corrections for the regular lattice.
IV Dynamics of the Pairwise Correlation Function
We now formulate generalized dynamics for the distance-one pairwise correlation function, $F_{II}^{(1)}\left(t\right)$, for both the lattice and the RRG, using a series of hierarchical ODEs for equivalent correlation functions at a range of distances on the network, $\{F_{II}^{(r)}\left(t\right)\}^{r=r_{\text{max}}}_{r=1}$. On the RRG, which lacks a spatial embedding, the notion of Euclidean distance is ill-defined. As such, it is often preferable to consider the shortest path between nodes as a metric of distance. This notion of distance affords the definition of “shells” of nodes surrounding each node on a network, where shells of distance one contain all nodes directly connected to a given node, shells of distance two contain all nodes connected to the nodes of distance one that are not in shells of distance one, etc. (see Fig. 4). This, in turn, means that a set of correlation functions $\{F_{II}^{(\ell)}\left(t\right)\}^{\ell=\ell_{\text{max}}}_{\ell=1}$ can be defined in a manner of Eq. (4), where $\ell$ denotes the shell. In Eqs. (4)-(5), the Euclidean distance $r$ appears only in the enumeration of the correlation function and is not explicitly involved in the system dynamics. As such, Euclidean distance can be exchanged with the shell number $\ell$, where every possible distance from the origin in the lattice is exchanged by the corresponding shell number $\ell$. For example, for the two-dimensional lattice, the distances $r=1,\sqrt{2}/2,\sqrt{3}/2,\dots$, may be exchanged with $\ell=1,2,3,\dots$. In the following sections, we proceed in terms of this notion of shell distance, rather than Euclidean distance.
The dynamics of the pairwise correlation function $F_{II}^{(\ell)}(t)=x_{II}^{(\ell)}/x_{I}^{2}$ can be derived by analyzing the change in both the two-point probability $x_{II}^{(\ell)}$ and the global infected density $x_{I}$. The dynamics of $x_{II}^{(\ell)}$ includes contributions from infection events in which a susceptible node in an IISS pair becomes infected through a third, infected neighbor distinct from the original $I$ node. This process is described by the three–point function $x_{ISI}^{(\ell\ell^{\prime}m)}$, which we approximate using the Kirkwood superposition approximation (KSA). The KSA expresses $x_{ISI}^{(\ell\ell^{\prime}m)}$ as a product of pairwise distributions,
$x_{ISI}^{(\ell\ell^{\prime}m)}\!=\!\frac{x_{SI}^{(\ell)}x_{SI}^{(m)}x_{II}^{(\ell^{\prime})}}{(1\!-\!x_{I})x_{I}^{2}}\!=\!(1-x_{I})x_{I}^{2}F_{SI}^{(\ell)}(t)F_{SI}^{(m)}(t)F_{II}^{(\ell^{\prime})}(t),$ (7)
where $\ell$, $\ell^{\prime}$, and $m$ are the shell distances between nodes 1-2, 1-3, and 2-3 of the triplet, respectively. Since infection occurs by direct contact between a susceptible node and a third infected neighbor, we set $m=1$ and sum over all possible third-party shells $\ell^{\prime}$, weighted by $p(\ell^{\prime}|\ell)$—the conditional probability that a neighbor of a susceptible node in shell $\ell$ lies in shell $\ell^{\prime}$. The total contribution is then equal to $2\beta(1-x_{I})F_{SI}^{(\ell)}(t)\sum_{\ell^{\prime}>0}p(\ell^{\prime}|\ell)\,F_{II}^{(\ell^{\prime})}(t)$, where the factor of two accounts for symmetric SSII and IISS configurations.
Another contribution to the dynamics of $x_{II}^{(\ell)}$ arises from the direct infection of adjacent IISS pairs, which converts them into IIII pairs. This occurs only when $\ell=1$, and can therefore be written as $2\beta\delta_{\ell,1}x_{IS}/x_{I}^{2}=2\beta\delta_{\ell,1}\left(1-x_{I}\right)F_{SI}^{(1)}\left(t\right)/x_{I}$, where again the factor of two follows similarly as before, and $\delta_{n,m}$ is the Kronecker delta. Furthermore, $F_{II}^{(\ell)}(t)$ is also affected by variations in the overall infection density $x_{I}$ that appears in the denominator of $F_{II}^{(\ell)}(t)$. Differentiating $F_{II}^{(\ell)}(t)=x_{II}^{(\ell)}/x_{I}^{2}$ yields a term $2x_{II}^{(\ell)}\dot{x}_{I}/x_{I}^{3}$, which equals $2F_{II}^{(\ell)}(t)\dot{x}_{I}/x_{I}$. Substituting $\dot{x}_{I}$ from Eq. (6) and using the identity $(1-x_{I})F_{SI}^{(\ell)}(t)=1-x_{I}F_{II}^{(\ell)}(t)$ as explained in Eq. (5) recasts this contribution as $2F_{II}^{(\ell)}(t)\,\beta k_{0}(1-x_{I})F_{SI}^{(1)}(t)$. Here, the factor of two reflects the squared normalization in $x_{I}$. Finally, although recovery events reduce both $x_{II}^{(\ell)}$ and $x_{I}$, their effects cancel out in the evolution of $F_{II}^{(\ell)}\left(t\right)$. Recovery of one node in an IIII pair contributes $2F_{II}^{(\ell)}\left(t\right)$ due to pair symmetry, while the change in $x_{I}$ from the same recovery contributes $x_{I}$ to $\dot{x}_{I}$, such that it modifies the denominator in a way that precisely offsets that reduction, resulting in a net zero contribution.
Putting it all together, the dynamics of $F_{II}^{(\ell)}$ are described by the coupled system of ODEs
$\frac{\mathrm{d}F_{II}^{(\ell)}(t)}{\mathrm{d}t}$ $=\left[F_{SI}^{(\ell)}(t)\sum_{\ell^{\prime}>0}p(\ell^{\prime}|\ell)F_{II}^{(\ell^{\prime})}(t)\!+\!\frac{\delta_{\ell,1}}{k_{0}x_{I}}\!-\!F_{II}^{(\ell)}(t)\right]$
[…]
where $\ell\in\{1,\dots,\ell_{\max}\}$, and $p(m|n)$ is the conditional probability that a given neighbor of a node in shell $n$ lies in shell $m$. The dynamics can also be written entirely in terms of $F_{II}^{(\ell)}$ by using Eq. (5), which is convenient for numerical solutions.
Notably, Eq. (8) applies both to the regular lattice and to the RRG. For example, in a two‑dimensional square lattice (Fig. 4(a)), a node in $(1,0)$ (first shell) has one neighbor in the zeroth shell, two in the second shell, and one in the third shell, such that $p(k|1)=\delta_{k,0}/4+\delta_{k,2}/2+\delta_{k,3}/4$. Similarly, a node in (1,1)(-1,1) (second shell) has two neighbors in the first shell and two in the fourth shell, giving $p(k|2)=\delta_{k,1}/2+\delta_{k,4}/2$. By translational invariance, the probabilities $p(\ell^{\prime}|\ell)$ are uniform across nodes, and hence are global constants that vary with the lattice dimension.
Figure 4: (a) A two-dimensional regular lattice with its origin at $(0,0)$. The purple, green, blue, and red circles represent the first, second, third, and fourth shells, respectively. (b) RRG shells with node degree $k_{0}=4$. The central node is black, the first shell nodes are colored purple, while the second shell nodes are colored green.
V Computing Shell Probabilities on Regular Random Graphs
To complete the formulation of the dynamics of the pairwise correlation function, Eq. (8), it is left for us to find an explicit expression for $p(\ell^{\prime}|\ell)$ in the case of the RRG. For the latter, the fraction of individuals located in shells beyond distance $\ell$ follows the Gompertz distribution [18]
$P\left(L>\ell\right)=\exp\left\{-\frac{k_{0}}{\left(k_{0}-2\right)N}\left(\left(k_{0}-1\right)^{\ell}-1\right)\right\},$ (9)
where $P\left(L>\ell\right)$ is the probability that a randomly selected node on the network lies at a greater distance than $\ell$ from a given node. The difference between the complementary cumulative fractions of consecutive shells, therefore, determines the fraction of nodes in a specific shell $p\left(\ell\right)=P\left(L>\ell-1\right)-P\left(L>\ell\right)$, and the number of nodes in the $\ell$-th shell is thus given by
$N_{\ell}=Np\left(\ell\right)=N\left[P\left(L>\ell-1\right)-P\left(L>\ell\right)\right].$ (10)
The conditional probability $p(k|n)$ quantifies the likelihood that a randomly chosen neighbor of a node in the $n$-th shell belongs to the $k$-th shell, and plays a crucial role in characterizing network connectivity at increasing distances from a reference node. At this point, we note that the number of nodes in the $\ell$-th shell satisfies
$N_{\ell}=\prod_{k=0}^{\ell-1}\frac{p(k+1\,|\,k)}{p(k\,|\,k+1)},$ (11)
where this identity stems from Bayes’ theorem
$p(k\,|\,k+1)=p(k+1\,|\,k)\frac{p(k)}{p(k+1)}=p(k+1\,|\,k)\frac{N_{k}}{N_{k+1}},$ (12)
with $p(k)$ being proportional to the shell population $N_{k}$.
The calculation of the conditional probabilities $p(\ell-1|\ell)$, $p(\ell|\ell)$, and $p(\ell+1|\ell)$ for a target node located in the $\ell$-th shell, is performed iteratively. At each step, these probabilities are determined using the previously computed values for the $(\ell-1)$-th shell and the shell populations, as given by Eq. (10). Specifically, the recursion relation exploits the previously obtained $p(\ell|\ell-1)$, and all the shell populations up to the current $\ell$-th shell, $N_{0},\dots,N_{\ell}$ to advance to the next shell, ensuring that all dependencies are consistently propagated throughout the shell hierarchy. From Eq. (12), it follows that $p\left(\ell-1|\ell\right)=p\left(\ell|\ell-1\right)N_{\ell-1}/N_{\ell}$. Now, the probability $p(\ell|\ell)$ describes the likelihood that a neighbor of a node in shell $\ell$ also lies in the same shell. To compute this, we first condition on the alternative—that the edge does not point back to shell $\ell-1$—which provides the prefactor $\left(1-p(\ell-1|\ell)\right)$. The remaining neighbors of the node can then be distributed either within the same shell $\ell$ or into the next shell $\ell+1$. To evaluate this balance, we count the total number of available “stubs” in shell $\ell$ that can remain internal $k_{0}N_{\ell}$, subtracting those already used to connect back into previous shell $\ell-1$, $k_{0}p\left(\ell-1|\ell\right)N_{\ell}$, and further excluding the trivial self‑connection of the test node $k_{0}\left(1-p\left(\ell-1|\ell\right)\right)$. This yields the numerator of $p(\ell|\ell)$ to be $k_{0}N_{\ell}-k_{0}p\left(\ell-1|\ell\right)N_{\ell}-k_{0}\left(1-p\left(\ell-1|\ell\right)\right)$, reflecting the number of ways that edges can close within shell $\ell$ once the backward connections and self-loops have been removed. The denominator counts the total remaining available “stubs” in the network after all edges to shells $0,\dots,\ell\,$ have been allocated $k_{0}N-\sum_{k=0}^{\ell}k_{0}N_{k}$, plus the corrected contribution from current $\ell$-th shell. After some simplifications, the resulting expression reads
$p(\ell|\ell)=\frac{(N_{\ell}-1)\,\left[1-p(\ell-1|\ell)\right]^{2}}{N-\sum\limits_{k=0}^{\ell}N_{k}+(N_{\ell}-1)\left[1-p(\ell-1|\ell)\right]}.$ (13)
This expression captures the self‑closing probability within the shell as the ratio of “internal same‑shell connections” to “all possible remaining connections”.
The probability $p(\ell+1|\ell)$, representing the likelihood that a neighbor of a node in shell $\ell$ resides in the forward shell $\ell+1$, can be formulated either by enforcing the complementary probability, $p(\ell+1|\ell)=1-p(\ell|\ell)-p(\ell-1|\ell)$, or by employing the same combinatorial logic where the numerator corresponds to available endpoints in the forward shells and the denominator accounts by all remaining stubs. The system is rendered self-consistent by stipulating the boundary condition $p(1|0)=1$, ensuring that the initial shell is reached deterministically from the origin. With this assignment, the recursive relations for the shell populations $N_{k}$ and the transition probabilities $p(k|\ell)$ are fully formulated, thereby defining the complete hierarchy of shell-wise network topology.
To assess the accuracy of the pairwise correlation correction to mean-field theory on the RRG, we directly compare solutions using our analytical framework to numerical simulations of the SIS model. Specifically, we solve a system of $n$ coupled ODEs for the pairwise correlation functions in each shell $\ell$ as given by Eq. (8), with the number of shells $n$ naturally limited so that $\sum_{i=1}^{n}N_{i}\approx N$, and truncated at the shell where $p(n+1|n)<1/N$, indicating that there is fewer than one node in the subsequent shell. The conditional probabilities determine how correlations propagate through the network topology, ensuring consistency in the shell structure. Once solved, the hierarchy of correlation functions, particularly the distance-one correlation function, is then used in Eq. (6) to predict the global infection density. This unified approach enables direct comparison between theory and simulation. We illustrate this on a network of $10^{4}$ nodes, with an initial fraction $10^{-2}$ of infected individuals. The results shown in Fig. 2 show good agreement between the numerical solution of Eq. (8) (blue) and corresponding discrete simulations (red). We note a slight discrepancy between our corrections to mean-field theory and corresponding simulations for infection rates $\beta$ close to the critical values $\beta_{c}k_{0}/\gamma=1.33$ and $1.20$ for $k_{0}=4$ and $k_{0}=6$, respectively, according to the formula $\beta_{c}k_{0}/\gamma=k_{0}/(k_{0}-1)$ for homogeneous networks [19]. These critical values separate the regime where the infection persists in the population from the regime in which the epidemic dies out deterministically. The discrepancies near these critical points, where epidemic dynamics are dominated by strong fluctuations and long-range correlations, are inherently difficult to attenuate in low-order moment closure, such as the KSA, and have been discussed in previous works [11, 12]. In general, the agreement between model simulations and numerical solutions using our modified theory improves with the increase of the node degree $k_{0}$ or as the infection rate $\beta$ increases.
VI A modified steady state analysis on regular random networks.
Alongside the dynamical trajectory of infection density, we now compute the steady-state value of infection density, $\bar{x}_{I}$, within the corrected mean-field framework, which corresponds to the so-called endemic state. This is of particular interest in percolation theory, where the non-equilibrium phase transition between endemicity and elimination of the virus at a critical rate of infection, $\beta_{c}$, is not fully understood on a wide range of network topologies. Setting $\mathrm{d}x_{I}/\mathrm{d}t=0$ in Eq. (6), we find
$\bar{x}_{I}=\bar{x}_{I}^{\text{MF}}/\bar{F}_{II}^{(1)},$ (14)
where $\bar{x}_{I}^{\text{MF}}=1-\gamma/(\beta k_{0})\equiv 1-1/R_{0}$ is the standard mean-field prediction for the endemic state, obtained by finding the steady state of Eq. (6) with $F_{II}^{(1)}(t)=1$, while $\bar{F}_{II}^{(1)}$ is the steady-state value of the distance-one pairwise correlation function.
For a RRG, $\bar{F}_{II}^{(1)}$ is obtained by setting $\mathrm{d}F_{\ell}\left(t\right)/\mathrm{d}t=0$ in Eq. (8) and using the the expression for $\bar{x}_{I}^{\text{MF}}$, yielding
$\bar{F}_{II}^{(\ell)}=\frac{\bar{F}_{II}^{(1)}-\bar{F}_{II}^{(\ell)}\bar{x}_{I}^{\text{MF}}}{\bar{F}_{II}^{(1)}-\bar{x}_{I}^{\text{MF}}}\!\!\!\!\!\!\sum_{\ell^{\prime}=\max\{\ell-1,1\}}^{\ell+1}\!\!\!\!\!\!\!\!\!\!p\left(\ell^{\prime}|\ell\right)\bar{F}_{II}^{(\ell^{\prime})}+\frac{\bar{F}_{II}^{(1)}\delta_{\ell,1}}{k_{0}\bar{x}_{I}^{\text{MF}}}.$ (15)
Analysis of the steady-state behavior of the pairwise correlation function on the RRG shows an exponential decay of $\bar{F}_{II}^{(\ell)}$ with increasing distance between nodes (Figs. 5(a)-(b)). This decay is accurately described by
$\bar{F}_{II}^{(\ell)}=1+\alpha e^{-\lambda\ell}\equiv 1+\alpha K^{\ell},$ (16)
where $\alpha$ and $K=\exp(-\lambda)$ are parameters yet to be determined. Employing this empirical dependence allows finding $\bar{F}_{II}^{(\ell)}$. Substituting Eq. (16) into (15) for shells $\ell=1,2$ and further using Eq. (16) for $\ell=3$, yields two implicit equations for $\alpha$ and $K$. Following algebraic manipulations detailed in Appendix A, these equations reduce to a cubic equation $AK^{3}+BK^{2}+CK+D=0$, alongside an explicit equation connecting $\alpha$ and $K$, which can be solved numerically for any $k_{0}$. Notably, while Eq. (15) breaks down at larger values of $\ell$ (see Fig. 5(b)), it remains rather accurate for $\ell={\cal O}(1)$; this allows for an accurate approximation of the distance-one pairwise correlation, which is the primary quantity of interest in our analysis, as it determines the correction to $\bar{x}_{I}$.
Figure 5: (a)-(b) The steady-state pairwise correlation function, $\bar{F}_{II}^{\ell}$, versus the network distance $\ell$, on linear and semi-logarithmic scales, respectively. Circles indicate values obtained from numerical simulations, while solid lines represent exponential fits based on Eq. (16). (c)-(d) The steady-state pairwise correlation function $\bar{F}_{II}^{1}$ and steady-state infected fraction $\bar{x}_{I}$, respectively, versus the average node degree $k_{0}$. Numerical solutions (circles) and simulation values (triangles) show strong agreement. Additionally, power series approximations of the second order of $k_{0}^{-1}$ (solid green lines) converge toward the numerical and simulation results as $k_{0}$ increases.
The analysis above greatly simplifies under the assumption of $k_{0}\ll N^{1/2}$, making the network locally tree like  [^1] In the complementary regime $k_{0}\gtrsim N^{1/2}$, the regular random graph is no longer locally tree-like and finite-size effects as well as an increased prevalence of short cycles in the network occurs. However, since the expected number of shells in a RRG scales as $\log_{k_{0}}\!N$, for $k_{0}\gtrsim N^{1/2}$ the network has only $\sim 2$ shells, which is clearly a less interesting regime to deal with.. In this limit, the conditional probabilities read
$p(0|0)$ $=0,\quad p(1|0)=1,$
$p(0|1)$ $=1/k_{0},\quad p(1|1)=\mathcal{O}(k_{0}/N),$
$p(2|1)$ $=1-1/k_{0}+\mathcal{O}(k_{0}/N),$
$p(1|2)$ $=1/k_{0}+\mathcal{O}(k_{0}/N),\quad p(2|2)=\mathcal{O}(k_{0}^{2}/N),$
$p(3|2)$ $=1-1/k_{0}+\mathcal{O}(k_{0}^{2}/N).$ (17)
Substituting these results into Eq. (15) together with the ansatz from Eq. (16), we obtain
$\alpha=-\frac{1}{\left(k_{0}-1\right)\left(R_{0}-1\right)K\left(K-R_{0}\right)+R_{0}K},$ (18)
$0=K^{3}-R_{0}\frac{\left(2k_{0}-1\right)}{k_{0}}K^{2}+\left[R_{0}^{2}-\frac{R_{0}-1}{k_{0}-1}\right]K-\frac{R_{0}}{k_{0}}.$ (19)
Further assuming $k_{0}\gg 1$, and expanding $K$ and $\alpha$ in powers of $k_{0}^{-1}$ in the regime $1\ll k_{0}\ll N^{1/2}$ gives
$K$ $=\frac{1}{R_{0}k_{0}}+\frac{R_{0}+1}{R_{0}^{3}k_{0}^{2}}+\mathcal{O}\left(\frac{1}{k_{0}^{3}}\right),$ (20)
$\alpha$ $=\frac{1}{R_{0}-1}+\frac{R_{0}^{2}-R_{0}+1}{R_{0}(R_{0}-1)^{2}k_{0}}+\mathcal{O}\left(\frac{1}{k_{0}^{2}}\right).$ (21)
Here, we have retained terms in the series expansion of $\alpha$ up to ${\cal O}(k_{0}^{-1})$, since the leading term in the expansion of $K$ is also of ${\cal O}(k_{0}^{-1})$, while we are interested in expressing their product $\bar{F}_{II}^{(1)}=1+\alpha K$ up to ${\cal O}(k_{0}^{-2})$. This yields
$\bar{F}_{II}^{(1)}\!=\!1\!+\!\frac{1}{(R_{0}\!-\!1)R_{0}k_{0}}\!+\!\frac{R_{0}^{3}+R_{0}-1}{(R_{0}\!-\!1)^{2}R_{0}^{3}k_{0}^{2}}\!+\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\!.$ (22)
Using (14), this yields the steady-state infection density
$\bar{x}_{I}=\bar{x}_{I}^{\text{MF}}\left[1\!-\!\frac{1}{R_{0}\left(R_{0}\!-\!1\right)k_{0}}\!-\!\frac{R_{0}^{3}-1}{R_{0}^{3}\left(R_{0}\!-\!1\right)^{2}\!k_{0}^{2}}\!-\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\!\right]\!\!.$ (23)
Comparisons between our predictions for $\bar{F}_{II}^{(1)}$ and $\bar{x}_{I}$ and numerical simulations are shown in Figs. 5(c)-(d) and excellent agreement is observed at large $k_{0}$. Notably, these expansions make explicit how finite-degree corrections alter both the two-point correlation function and infection density relative to the standard mean-field predictions. The consistent negativity of the correction terms in $\bar{x}_{I}$ reflects the dampening effect of local clustering and correlations on disease spread in finite-connectivity networks, consistent with the fact that $\bar{F}_{II}^{(1)}(t)\geq 1$ at all distances, as seen in Fig. 2.
Finally, exploiting the power series expansions of the steady-state infection density $\bar{x}_{I}$ and pairwise correlation function $\bar{F}_{II}^{(1)}$, we also performed a linear stability analysis of the coupled epidemic dynamics given by Eqs. (6) and (8) for $\ell=1$, assuming that the relaxation time of all $\bar{F}_{II}^{(\ell)}$ is comparable to that of $\bar{F}_{II}^{(1)}$, for $\ell>1$. By evaluating the Jacobian matrix at the endemic equilibrium, we explicitly capture corrections beyond the mean-field limit. Expanding the resulting relaxation time (corresponding to the lowest eigenvalue) up to ${\cal O}(k_{0}^{-2})$ we find
$\tau=\frac{1}{R_{0}\!-\!1}\!\!\left[1\!+\!\frac{4R_{0}-2}{\left(R_{0}\!-\!1\right)^{2}k_{0}}\!+\!\frac{2\left(4R_{0}\!-\!7\right)R_{0}^{3}\!+\!2}{\left(R_{0}\!-\!1\right)^{3}R_{0}^{2}k_{0}^{2}}\!+\!\mathcal{O}\!\!\left(\!\frac{1}{k_{0}^{3}}\!\right)\right]\!.$ (24)
The zeroth-order value $1/(R_{0}-1)$ is the mean-field relaxation time, while the higher-order positive corrections imply that for lower $k_{0}$, the system takes more time to converge.
VII Discussion
In this work, we have characterized, through numerical simulations and analytical calculations, the role of spatial correlations in the transmission dynamics of SIS processes across various network topologies. Our findings demonstrate that spatial correlations in infectiousness are suppressed by decreasing the average shortest path length by random edge rewiring and increasing the infection rate. The observed sensitivity of spatial correlations to network topology implies that corrections to mean-field theory previously derived for regular lattices are unsuitable for RRGs. To address this, we leveraged existing analytical results on RRG connectivity to develop appropriate mean-field corrections tailored specifically to regular random topologies. These new corrections provide significantly improved predictions of infection dynamics in SIS processes on RRGs, particularly for values of the basic reproduction number above the epidemic threshold (i.e., $R_{0}>1$). The transition at $R_{0}=1$ defines the critical epidemic threshold separating disease extinction from persistence. Traditionally, the mean-field approximation predicts a steady-state infected fraction $\bar{x}_{I}^{\text{MF}}=1-1/R_{0}$. However, our corrections adjust this estimate and reflect spatial correlations arising from clustering of infections. This adjustment reduces the predicted endemic prevalence, capturing the clustering effect of infected nodes. As a result, the effective pool of susceptible neighbors per infected individual is smaller than assumed in mean-field theory, which in turn lowers the steady-state infection density.
While the corrected mean-field approach outperforms previous lattice-based corrections in this regime, the accuracy of the predictions still depends on the proximity of the rate of infection, $\beta$, to the critical value for a given underlying network structure. Notably, the accuracy of these corrections diminishes near the critical threshold, indicating the need for more robust approximation methods in near-critical regimes.
Future research could attempt to re-derive corrections for the RRG by invoking the Fisher-Kopeliovich approximation, which is known to provide improved accuracy where the KSA becomes inadequate [21] (albeit at the expense of greater analytical complexity). Another potential direction would be to expand the current framework to topologies on which node degree is heterogeneous. This will enable us to predict and analyze more realistic scenarios, particularly where degree variability plays a crucial role in epidemic outcomes [22].
Overall, our work broadens the applicability of existing corrections to mean-field theory and highlights how network structure impacts spatial correlations in SIS processes. A promising direction for future research involves exploring intermediate network structures to bridge the analytical gap between regular lattices and fully randomized network topologies.
Appendix A Corrections to the steady state density
Here we derive closed-form equations for $\alpha$ and $K$ in the main text, which allow finding the correction to the steady-state infection density for arbitrary $k_{0}$. Writing Eq. (15) for shells $\ell=1,2$ yields
$\bar{F}_{II}^{(1)}\!=\!\frac{\bar{F}_{II}^{(1)}\!-\!\bar{F}_{II}^{(1)}\bar{x}_{I}^{\text{MF}}}{\bar{F}_{II}^{(1)}-\bar{x}_{I}^{\text{MF}}}\!\!\left[p\left(1|1\right)\bar{F}_{II}^{(1)}\!+\!p\left(2|1\right)\bar{F}_{II}^{(2)}\right]\!+\!\frac{\bar{F}_{II}^{(1)}}{k_{0}\bar{x}_{I}^{\text{MF}}},$
$\bar{F}_{II}^{(2)}\!=\!\frac{\bar{F}_{II}^{(1)}\!-\!\bar{F}_{II}^{(2)}\bar{x}_{I}^{\text{MF}}}{\bar{F}_{II}^{(1)}-\bar{x}_{I}^{\text{MF}}}\!\!\left[p\!\left(1|2\right)\!\bar{F}_{II}^{(1)}\!+\!p\!\left(2|2\right)\!\bar{F}_{II}^{(2)}\!+\!p\!\left(3|2\right)\!\bar{F}_{II}^{(3)}\!\right]\!.$
Substituting Eq. (16) for shells $\ell=1,2,3$ into these equations, and using the fact that 1=p(k+1|k)+p(k|k)+p(k1|k)1=p(k+1|k)+p(k|k)+p(k-1|k), we obtain two coupled equations for $\alpha$ and $K$:
$\alpha K=-\frac{1-k_{0}\bar{x}_{I}^{\text{MF}}p\left(0|1\right)}{\frac{1-k_{0}\bar{x}_{I}^{\text{MF}}}{1-\bar{x}_{I}^{\text{MF}}}+k_{0}\bar{x}_{I}^{\text{MF}}\left(p\left(1|1\right)+p\left(2|1\right)K\right)},$ (25)
and
$0=p\left(3|2\right)\bar{x}_{I}^{\text{MF}}K^{3}+\left(\bar{x}_{I}^{\text{MF}}p\left(2|2\right)-p\left(3|2\right)\right)K^{2}$ (26)
$\hphantom{=}+\left(p\left(1|2\right)\bar{x}_{I}^{\text{MF}}-p\left(2|2\right)+1\right)K-p\left(1|2\right)$
$\hphantom{=}-\frac{1-\bar{x}_{I}^{\text{MF}}}{\alpha K}\left[p\left(3|2\right)K^{2}\!+\!p\left(2|2\right)K\!+\!p\left(1|2\right)\right]+\frac{K\bar{x}_{I}^{\text{MF}}}{\alpha K}\!.$
Finally, substituting the value of $\alpha$ from Eq. (25) into Eq. (26) and rearranging terms yields a cubic equation for $K$, $AK^{3}+BK^{2}+CK+D=0$, where the coefficients satisfy:
$A$ $=k_{0}{x}_{I}^{\text{MF}2}p\left(3|2\right)\left(1-p\left(1|1\right)\right)$
$\hphantom{=}-k_{0}{x}_{I}^{\text{MF}}p\left(3|2\right)p\left(2|1\right)-p\left(3|2\right),$ (27)
$B$ $=k_{0}{x}_{I}^{\text{MF}2}\left[p\left(2|2\right)-p\left(1|1\right)p\left(2|2\right)+p\left(1|1\right)p\left(3|2\right)\right]$
$\hphantom{=}+k_{0}{x}_{I}^{\text{MF}}p\left(2|1\right)\left[1+p\left(3|2\right)-p\left(2|2\right)\right]-p\left(2|2\right){x}_{I}^{\text{MF}},$ (28)
$C$ $=k_{0}{x}_{I}^{\text{MF}2}\left[p\left(1|2\right)-p\left(1|1\right)p\left(1|2\right)+p\left(1|1\right)p\left(2|2\right)\right]$
$\hphantom{=}+k_{0}{x}_{I}^{\text{MF}}\left[1-p\left(2|1\right)\left(1-p\left(2|2\right)+p\left(1|2\right)\right))\right]$
$\hphantom{=}-p\left(1|2\right){x}_{I}^{\text{MF}}+\frac{{x}_{I}^{\text{MF}}\left(1-k_{0}\right)}{1-{x}_{I}^{\text{MF}}},$ (29)
$D$ $=k_{0}{x}_{I}^{\text{MF}2}p\left(1|1\right)p\left(1|2\right)+k_{0}{x}_{I}^{\text{MF}}p\left(2|1\right)p\left(1|2\right).$ (30)